40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Feb 28 12:00:54 2012, kiwamu, Update, LSC, installed anti-whitening filters 
    Reply  Tue Feb 28 15:48:32 2012, kiwamu, Update, LSC, installed anti-whitening filters 
    Reply  Tue Feb 28 16:39:25 2012, kiwamu, Update, LSC, MICH and PRCL signals in a simulation REFL11_PRMI.pngREFL55_PRMI.png
       Reply  Tue Feb 28 16:44:56 2012, rana, Update, LSC, MICH and PRCL signals in a simulation 
Message ID: 6330     Entry time: Tue Feb 28 12:00:54 2012     Reply to this: 6331   6334
Author: kiwamu 
Type: Update 
Category: LSC 
Subject: installed anti-whitening filters 

I found that none of the filter banks in the LSC input signals have the precise anti-whitening filters.

I installed the precise filters on REFL11, REFL33, REFL55 and AS55 based on Jenne's measurement (#4955)

After installing them I briefly checked the REFL11 sensing matrix with the PRMI locked, but it didn't change so much from what I got (#6283).

But I felt that the PRMI became more robust after that ... I just felt so ...

 


(Background)

The lock of the PRMI doesn't look healthy, especially the sensing matrix doesn't make sense at all (#6283).
A very staring thing in the sensing matrix is that the REFL11 and REFL55 didn't show the 90 degree separation between MICH and PRCL.
So I suspected some electronics, particularly the demodulation boards.
 

(What I did)

I checked the anti-whitening filters shape to see if they are ok or not.
I found that they all had the default filters of two zeros at 150 Hz and two poles at 15 Hz. So they weren't quite tuned.
I thought this could be a problem when I measure the sensing matrix because I usually excite the length DOFs at a high frequency of 283.1 Hz
and the mismatches between the anti-whitening and whitening filters may lead to something funny at such a high frequency.
 
So I installed the precise filters on REFL11, REFL33, REFL55 and AS55.
After that I did a orthogonality test on each I-Q pair of the demod signals to correct the D-phases and the relative gain between I and Q.

 

(Next ?)

 Rana and I discussed the plan and decided to go back to a simple Michelson which should be easy enough to understand what is going on and should allow us a complete set of measurements.
Our big concern behind it is that we maybe locking the PRMI at a funny operation point.
In order to assess the issue I will do the following actions on the Michelson at first and then apply the same things on the PRMI later :
  • Check  the amount of of the sidebands using the OSA
  • Check the amount of the DC light
  • Check the sensing matrix to see if the absolute values in watt / meter make sense or not
    • This work needs calibrations on all the demodulated board (this is equivalent to measuring the conversion losses of the mixers in the demod boards).
  • Measure the contribution from the RAMs (it must be measurable by some means)
ELOG V3.1.3-